# FissionVAE: Federated Non-IID Image Generation with Latent Space and Decoder Decomposition The source code of FissionVAE is publicly available at: https://github.com/Rand2AI/FissionVAE

### v1

@@ 5 Conclusion
In this study, we introduce FissionVAE, a generative model designed for federated image generation in non-IID data environments. By decomposing the latent space and employing decoder branches tailored to specific client groups, FissionVAE not only enhances generation quality but also maintains the distinct visual features of diverse data subsets. Our experiments utilize composite datasets includeing the Mixed MNIST and a broader dataset CHARM comprising cartoon and human faces, animals, marine vessels, and remote sensing images. Evaluation results demonstrate that FissionVAE significantly outperforms baseline federated VAE models in terms of image quality. Notably, the implementation of heterogeneous decoder branches and the adoption of wave encoding for defining priors have proven effective, underscoring our model’s capability to adapt to the varied demands of non-IID data distributions within federated learning frameworks.
Several avenues appear promising for further research and improvement of FissionVAE. Enhancing the stability and optimization dynamics of heterogeneous decoder branches could address challenges encountered during training and generation, particularly as the complexity of data and model architecture increases. Exploring the potential of FissionVAE to support cross-modality data generation could pave the way for its application in generating vision-language data within a federated environment. Moreover, as real-world federated learning scenarios often involve an increasing number of client groups, developing more sophisticated strategies to scale latent space and decoder decomposition will be crucial.

### v2

@@ 4 Conclusion
We presented FissionVAE, a generative model for federated image generation in non-IID data settings. By decoupling the latent space and employing group-specific decoder branches, FissionVAE enhances generation quality while preserving the distinct features of diverse data subsets. Experiments on Mixed MNIST and CHARM datasets demonstrated significant improvements over baseline federated VAE models, with heterogeneous decoder branches and wave-encoded priors proving particularly effective.
Future work includes improving the stability of heterogeneous decoder branches, enabling cross-modality data generation, and developing scalable strategies for handling an increasing number of client groups in real-world federated learning scenarios.